OpenStudy (anonymous):

~*~*~*~HELP~*~*~*~ A tuition scale for a state university shows an increase of $1,500 each year. Part 1: If the initial cost was $23,000, what will the cost of tuition be in twenty five years? Part 2: In complete sentences, explain how an arithmetic sequence was applied.

OpenStudy (anonymous):

Part #1: It will be 23,000 + 1500 times 25 Part #2: Each year, the tuition will be higher by $1,500. So, the year-over-year difference is "fixed" at $1,500. Therefore, tuition for various years from now forms an Arithmetic sequence.

OpenStudy (anonymous):

It said I had to use a different formula

OpenStudy (anonymous):

idk what formula tho

OpenStudy (anonymous):

@KingGeorge

OpenStudy (anonymous):

wait nevermind the formula looks like: an = a1 + (n-1)d

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

@Zarby @zainab_46 @sandbeach @radar @Meriam

OpenStudy (anonymous):

don't u times 1500 by 25?

OpenStudy (austinl):

\(\large{P_{Tuition} = 23,000 + 1,500t}\) Where \(\large{t=Number~of~Years}\)

OpenStudy (anonymous):

Ohk thank you all!

OpenStudy (anonymous):

still need help?

OpenStudy (anonymous):

no I think I'm good that was my last question but thanks(:

OpenStudy (anonymous):

no worries (: